Fundraising- The St. Marks Community Barbecue served 250 dinners. A child's plate cost 3.50 and an adult's plate cost 7.00. A total of 1347.50 was collected. How many of each type of plate was served?
Answer by checkley77(12844)

a+c=250 or a=250-c
7a+3.5c=1,347.50
7(250-c)+3.5c=1,347.50
1,750-7c+3.5c=1,347.50
-3.5c=1,347.5-1,750
-3.5c=-402.5
c=-402.50/-3.5
c=115 is the number of children plates.
a=250-115=135 is the number of adult plates sold.
Proof
7*135+3.5*115=1,374.50
945+402.5=1,374.50
1,374.50=1,374.50